• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name S. Strauch

    Log entry time 16:03:01 on October 29,1999

    Entry number 29288

    keyword=Day Shift Summary

    Day Shift Summary

    08:00 continued H-arm cycle with gamma-d after correcting the E-arm angle.
    08:34 - 10:13 : gamma-d runs (2731 - 2733), total Charge = 0.26 C.
    10:16 started e-d runs, first with 30 uA, later with 40 uA beam current
    10:16 - 11:35 : e-d runs (2734 - 2735), total Charge = 0.156 C
    11:35 - target change (2H to 1H)
    11:41 - 11:51 : e-p run (2736), total Charge = 0.023 C
    11:55 - 12:48 : gamma-p run (2737), total Charge = 0.090 C
    12:54 - 13:04 : 12C run
    13:05 : start Moeller setup, measurement, ...
    13:17 - 13:26 : 1000 Hz Pulser run (2738) PS1-5 = 65000, PS8 = 1
    14:00 : Problems to move electron spectrometer. Need an hall access later, tech on call is informed. Access planned during system checks.
    14:10 : VDC HV E_TOP and E_BTM was down. Resetted the VDC control and went back to 4.0 kV.
    16:00 : Moeller measurement is done, MCC is tuning back for exp. settings.